# K-Space-Aware Cross-Modality Score for Synthesized Neuroimage Quality Assessment

### v1

@@ III-C2 Similarity Loss
III-C3 Tumor Loss
The tumor loss function consists of a Laplacian loss function $\mathcal{L}_{lap}$ and the LPIPS loss function $\mathcal{L}_{lpips}$ [41]. The Laplacian loss function is defined as :
@@ III-C3 Tumor Loss
In (15), $\phi(\cdot)$ represents the feature extractor and $\tau(\cdot)$ computes the feature score from the k-th layer of the backbone architecture. As a result, the LPIPS value is the average score of all backbone layers. To compute the LPIPS loss, we used a well-trained VGG [30] network. The Laplacian loss is used to identify the tumor region’s high-frequency component. Due to LPIPS loss, the real tumor region and the reconstructed tumor region are more similar, which is more consistent with the radiologist’s judgment.
III-C4 Structure Loss
@@ III-C6 Total Loss
TABLE I: Description for notations used in our algorithms.
@@
Input : $s$, $t$, $\varphi_{E}$, $\varphi_{D}$, $\psi_{E}$, $\psi_{D}$, $\xi_{E}$, $\xi_{D}$, $\nu$, $FT$, $l$, $\zeta$, and $z$
Output : $\theta^{1}_{\varphi_{E}}$, $\theta_{\varphi_{D}}$, $\theta^{1}_{\psi_{E}}$, $\theta_{\psi_{D}}$, $\theta^{1}_{\xi_{E}}$, and $\theta_{\xi_{D}}$
@@ 4 $z_{t}\leftarrow\xi_{E}(t)$ // convert structure feature of target neuroimage
Input : $\hat{t}$, $\varphi_{E}$, $\theta^{1}_{\varphi_{E}}$, $\psi_{E}$, $\theta^{1}_{\psi_{E}}$, $\xi_{E}$, $\theta^{1}_{\xi_{E}}$, $n_{c}$, and $n_{nat}$
Output : $\theta_{n_{c}}$, $\theta_{n_{nat}}$
@@ 9 Ranking all $\eta_{total}$ and get in line with $\eta_{ra}$
Input : $\hat{t}$, $\varphi_{E}$, $\theta^{1}_{\varphi_{E}}$, $\psi_{E}$, $\theta^{1}_{\psi_{E}}$, $\xi_{E}$, $\theta^{1}_{\xi_{E}}$, $n_{c}$, $\theta_{n_{c}}$, $n_{nat}$, $\theta_{n_{nat}}$, $FT$, and $\nu$
Output : $\eta_{total}$
@@ 2 Initialize: start = 0, pairwise $\leftarrow\phi$, uniform_result $\leftarrow\phi$
III-D Algorithms
The two stages of training K-CROSS are depicted in Fig 2. The details of two-stage training algorithms and the inference algorithm are described in Algorithm 1, Algorithm 2 and Algorithm 3, respectively. For clarity, Table I provides notation descriptions that occurred in our algorithms.
IV NIRPS Dataset and Radiologist Score
To comprehensively evaluate the synthesis performance, we construct a large-scale multi-modal neuroimaging perceptual similarity (NIRPS) dataset with 6,000 radiologist judgments. NIRPS dataset is composed of three subsets generated by CycleGAN [44], MUNIT [14] and UNIT [21]. Each set contains 800 images generated by IXI and 1,200 images generated by BraTS. The IXI dataset includes two modalities, PD and T2, while the BraTS dataset includes three modalities, T1, T2, and FLAIR. In both the IXI and BraTS datasets, we randomly select 10 slices for training and collect the training results after each epoch of the model trained over 40 epochs.
IXI [2] collects nearly 600 MR images from normal and healthy subjects at three hospitals. The MR image acquisition protocol for each subject includes T1, T2, PD-weighted images (PD), MRA images, and Diffusion-weighted images. In this paper, we only use T1 (581 cases), T2 (578 cases) and PD (578 cases) data to conduct our experiments, and select the paired data with the same ID from the three modes. The image has a non-uniform length on the z-axis with the size of $256$ on the x-axis and y-axis. The IXI dataset is not divided into a training set and a test set. Therefore, we randomly split the whole data as the training set (80%) and the test set (20%).
BraTS2021 [29, 3] is designed for brain disease analysis and diagnosis. The dataset of multi-institutional and pre-operative MRI sequences is made publicly available, and it includes both training data (1251 cases) and validation data (219 cases). Each 3D volume is 155$\times$240$\times$240 in size and is imaged by four sequences: T1, T2, T1ce, and FLAIR.
Training Data Processing To ensure data validity and diversity, we remove their skulls for each slice, by splitting the three-dimensional volume and choosing slices ranging from 50 to 80 on the z-axis. All images are cropped to $256$ pixels in size. During the training stage, we choose a total of 10k images from the IXI and BraTS2021 datasets.
@@ IV NIRPS Dataset and Radiologist Score
IV-1 Radiologist Score
The NIRPS dataset contains radiologist scores (RS) resulting from manual annotation for each image. It is worth noting that the radiologist score $RS$ includes 10 levels, i.e., $RS\in[0,0.1,0.2,..,0.9]$. The higher RS value indicates better-synthesized neuroimage quality. The radiologists give scores in accordance with the level of diagnosis and therapy by using the synthesized neuroimage. Fig. 8 gives the distribution result of $RS$. We can see that synthesized performance varies among the three models and the average $RS$ is in the middle.
IV-2 How Radiologists Assess?
@@
IV-3 How Radiologists Combine Their Evaluations?
We hire 10 radiologists to evaluate the quality of each synthesized neuroimage. We remove the highest score and the lowest score from all radiologists. Then the final score is averaged by the rest score from 8 radiologists.
@@ IV-3 How Radiologists Combine Their Evaluations?
V Experiment and Ablation Study
TABLE II: Performance comparison of various IQA methods on NIRPS. NIRPS includes three subsets generated by CycleGAN, MUNIT, and UNIT. Specifically, each set consists of 800 images for IXI and 1,200 images for BraTS. Here, IXI and BraTS training sets are used to train our model. The inconsistency is computed, where $\downarrow$ is better. The best result is highlighted in red.
@@ V Experiment and Ablation Study
MAE [7] 0.3127 0.3335 0.1807 0.2368 0.2462 0.2462
PSNR [15] 0.3387 0.2060 0.1832 0.2515 0.2387 0.2585
SSIM [35] 0.3333 0.1998 0.1957 0.2437 0.2523 0.2645
MS-SSIM [36] 0.3245 0.1973 0.1934 0.2425 0.2487 0.2521
NLPD [19] 0.3128 0.1961 0.1921 0.2408 0.2466 0.2445
GMSD [39] 0.3069 0.1952 0.1910 0.2395 0.2421 0.2432
DeepIQA [4] 0.3052 0.1943 0.1827 0.2380 0.2402 0.2410
LPIP [41] 0.3023 0.1921 0.1786 0.2285 0.2398 0.2327
DIST [11] 0.3012 0.1901 0.1725 0.2291 0.2387 0.2230
@@
V-A K-CROSS vs Other Metrics
Table II illustrates the inconsistency between metrics and human evaluations of several datasets and generative models, with the highest performance shown in red. The calculation method for inconsistency value is given in Algorithm 4. We evaluate K-CROSS on datasets created by CycleGAN, MUNIT, and UNIT. The first column indicates various IQA methods. The second column indicates which datasets were used to train the K-CROSS model, including IXI or BraTS. From Table II, our proposed K-CROSS is more compatible with the assessments of radiologists. Note that the IXI dataset is a healthy person dataset. There is no lesion for each neuroimage. So K-CROSS only use the tumor branch and complex branch to assess the quality of neuroimage. The details are described in Section V-C.
V-B K-Space Importance
Table III records the ablation study of individual branches (the complex branch, tumor branch and structure branch) on various datasets. For instance, when we conduct the ablation study of the complex branch, K-CROSS remove the tumor branch and structure branch in the inference phase. In other words, K-CROSS only obtain $\eta_{complex}$ score. It applies the same setting for the other branches. It can be clearly observed that the complex branch obtain the highest score among the three branches. It strongly indicates the importance of k-space, which reflects the inherent properties of magnetic resonance imaging principles. The second best is the tumor branch. It also verifies the effectiveness of the tumor branch for the lesion disease dataset.
TABLE III: Effect of k-space importance (complex branch) and tumor branch. The inconsistency is computed, where $\downarrow$ is better. The best result is in red and the second best result is in blue.
@@ V-B K-Space Importance
V-C Metrics for Healthy Person
Table IV shows K-CROSS performance that surpasses the mainstream IQA methods on the IXI healthy-person dataset. As for assessing the synthesised neuroimage of healthy persons, K-CROSS removes the tumor branch in the inference phase. Because there are no lesions on healthy person datasets. It means that K-CROSS only combines $\eta_{complex}$ and the score of the structure encoder as the final score. From Table IV, it can be obviously observed that K-CROSS complex branch score $\eta_{complex}$ (blue value) has surpassed the other IQA methods, which identify the importance of k-space for MRI of healthy persons. Thus, K-CROSS still be able to serve as the metric for the synthesized quality of healthy person’s neuroimage.
TABLE IV: Metric performance for healthy persons on IXI. The inconsistency is computed, where $\downarrow$ is better. The best result is in red and the second best result is in blue.
@@ V-C Metrics for Healthy Person
MAE [7] 0.3160 0.1809 0.2442
PSNR [15] 0.3388 0.1835 0.2419
SSIM [35] 0.3352 0.1962 0.2423
MS-SSIM [36] 0.3238 0.1945 0.2418
NLPD [19] 0.3135 0.1918 0.2409
GMSD [39] 0.3090 0.1908 0.2398
DeepIQA [4] 0.3043 0.1856 0.2381
LPIP [41] 0.3022 0.1810 0.2374
DIST [11] 0.3056 0.1805 0.2356
@@
V-D Segmentation Network Effect
Table V shows K-CROSS remains stable performance even using different state-of-the-art medical segmentation models. Note that the parameters of the pre-trained segmentation network are frozen during the training phase. The first column denotes the segmentation method. We calculate the variance score of the K-CROSS value for CycleGAN, MUNIT and UNIT by using different segmentation backbone models. We find that the variance of K-CROSS performance is tiny (0.2%, 0.3%, and 0.2%). Hence, the performance of K-CROSS is not affected by the segmentation model.
TABLE V: The effect of segmentation network for K-CROSS value for various generative models of NIRPS.
@@ V-D Segmentation Network Effect
nnUnet [16] 0.1872 0.2246 0.2160
AttnUnet [28] 0.1873 0.2268 0.2157
SETR [43] 0.1852 0.2254 0.2162
CoTr [37] 0.1856 0.2244 0.2154
TransUNet [9] 0.1862 0.2247 0.2162
SwinUNet [6] 0.1845 0.2252 0.2158
TABLE VI: K-CROSS on NIRPS (CycleGAN). The inconsistency is computed, where $\downarrow$ is better. The best result is in red and the second best result is in blue.
@@
MAE [7] 0.3208 0.3321 0.3228
PSNR [15] 0.3190 0.3270 0.3261
SSIM [35] 0.3231 0.3231 0.3212
MS-SSIM [36] 0.3189 0.3180 0.3128
NLPD [19] 0.3176 0.3163 0.3090
GMSD [39] 0.3067 0.3154 0.3067
DeepIQA [4] 0.3058 0.3124 0.3045
LPIP [41] 0.3033 0.3056 0.3039
DIST [11] 0.3021 0.3032 0.3006
@@
TABLE VII: K-CROSS on the NIRPS dataset (MUNIT). The inconsistency is computed, where $\downarrow$ is better. The best result is in red and the second best result is in blue.
@@
MAE [7] 0.1872 0.1865 0.1842
PSNR [15] 0.1853 0.1873 0.1853
SSIM [35] 0.1967 0.1986 0.1982
MS-SSIM [36] 0.1932 0.1952 0.1923
NLPD [19] 0.1945 0.1943 0.1921
GMSD [39] 0.1920 0.1934 0.1932
DeepIQA [4] 0.1835 0.1845 0.1856
LPIP [41] 0.1783 0.1754 0.1784
DIST [11] 0.1742 0.1731 0.1786
@@
TABLE VIII: K-CROSS on the NIRPS dataset (UNIT). The inconsistency is computed, where $\downarrow$ is better. The best result is in red and the second best result is in blue.
@@
MAE [7] 0.2468 0.2445 0.2480
PSNR [15] 0.2421 0.2410 0.2410
SSIM [35] 0.2545 0.2505 0.2470
MS-SSIM [36] 0.2486 0.2474 0.2463
NLPD [19] 0.2481 0.2461 0.2450
GMSD [39] 0.2410 0.2388 0.2478
DeepIQA [4] 0.2398 0.2376 0.2391
LPIP [41] 0.2352 0.2341 0.2384
DIST [11] 0.2348 0.2337 0.2386
@@
V-E General Metric? Overcoming Domain Gap
The purpose of this paper is to demonstrate that K-CROSS is capable of serving as the standard measure for MRI datasets. We conduct extensive experiments and the results are given in Table VI, Table VII, and Table VIII to demonstrate that K-CROSS is not affected by dataset domain gap and the generative model. The training dataset is the BraTS dataset, and the test dataset is IXI. As described in Section V-C, we remove the tumor branch score, when K-CROSS evaluates the quality of neuroimage in healthy cases. We also observe that K-CROSS averagely surpasses DIST and LPIP (SOTA for natural image) by 7.8% and 16.5%, respectively, which proves that K-CROSS is built upon the basis of MRI principle instead of only on the natural image level. From this ablation study, we demonstrate that the performance of K-CROSS ($\mathcal{L}_{stru}$+$\mathcal{L}_{freq}$) is stable across several MRI datasets, with the potential to serve as a generic measure for evaluating the quality of the synthesized MRI.
VI Conclusion
In this paper, we proposed a new metric K-CROSS for assessing the performance of the synthesized medical images, which is built on the magnetic resonance imaging principle. To improve the capability of reconstruction during training K-CROSS, a complex U-Net was developed. As for training a learning-based full IQA metric, we further constructed a large-scale multi-modal neuroimaging perceptual similarity (NIRPS) dataset. Experimental results indicate that K-CROSS is a useful indicator for evaluating the quality of the generated medical data.
Limitation and Negative Society Impact Our method heavily relies on deep learning-based techniques but without directly injecting the knowledge of radiologists into K-CROSS. In the future, K-CROSS need to combine causal inference methods to enhance interpretability.

### v2

@@ 3.3.2 Similarity Loss
3.3.3 Tumor Loss
The tumor loss function consists of a Laplacian loss function $\mathcal{L}_{lap}$ and the LPIPS loss function $\mathcal{L}_{lpips}$ (Zhang et al., 2018). The Laplacian loss function is defined as :
@@ 3.3.3 Tumor Loss
In (15), $\phi(\cdot)$ represents the feature extractor and $\tau(\cdot)$ computes the feature score from the k-th layer of the backbone architecture. As a result, the LPIPS value is the average score of all backbone layers. To compute the LPIPS loss, we used a well-trained VGG (Simonyan and Zisserman, 2015) network. The Laplacian loss is used to identify the tumor region’s high-frequency component. Due to LPIPS loss, the real tumor region and the reconstructed tumor region are more similar, which is more consistent with the radiologist’s judgment.
3.3.4 Structure Loss
@@ 3.3.6 Total Loss
Table 1: Description for notations used in our algorithms.
@@
Input : Source neuroimage $s$, paired target neuroimage $t$, tumor encoder $\varphi_{E}$, source tumor decoder $\varphi_{D}$, source complex encoder $\psi_{E}$, complex decoder $\psi_{D}$, shared structure encoder $\xi_{E}$, shared structure decoder $\xi_{D}$, segmentation network $\nu$, Fourier transform $FT$, lesion mask $l$, k-space $\zeta$, structural feature $z$.
Output : The parameters of tumor encoder $\theta^{1}_{\varphi_{E}}$, tumor decoder $\theta_{\varphi_{D}}$, complex encoder $\theta^{1}_{\psi_{E}}$, complex decoder $\theta_{\psi_{D}}$, structure encoder $\theta^{1}_{\xi_{E}}$ and structure decoder $\theta_{\xi_{D}}$.
@@ 4 $z_{t}\leftarrow\xi_{E}(t)$ // convert structure feature of target neuroimage
Input : Synthesis target modality neuroimage $\hat{t}$, tumor encoder and its parameters from stage 1: $\varphi_{E}$ and $\theta^{1}_{\varphi_{E}}$, complex encoder and its parameters from stage 1: $\psi_{E}$ and $\theta^{1}_{\psi_{E}}$, structure encoder and its parameters from stage 1: $\xi_{E}$ and $\theta^{1}_{\xi_{E}}$, complex score network $n_{c}$, natural score network $n_{nat}$.
Output : The parameters of tumor encoder $\theta^{2}_{\varphi_{E}}$, complex encoder $\theta^{2}_{\psi_{E}}$, structure encoder $\theta^{2}_{\xi_{E}}$, complex score network $\theta_{n_{c}}$, natural score network $\theta_{n_{nat}}$
@@ 9 Ranking all $\eta_{total}$ and get in line with $\eta_{ra}$
Input : Synthesis target modality neuroimage $\hat{t}$, tumor encoder and its parameters from stage 1: $\varphi_{E}$ and $\theta^{1}_{\varphi_{E}}$, complex encoder and its parameters from stage 1: $\psi_{E}$ and $\theta^{1}_{\psi_{E}}$, structure encoder and its parameters from stage 1: $\xi_{E}$ and $\theta^{1}_{\xi_{E}}$, complex score network and its parameters from stage 2: $n_{c}$ and $\theta_{n_{c}}$, natural score network and its parameters from stage 2: $n_{nat}$ and $\theta_{n_{nat}}$, Fourier transform $FT$, segmentation network $\nu$.
Output : K-CROSS score $\eta_{total}$.
@@ 2 Initialize: start = 0, pairwise $\leftarrow\phi$, uniform_result $\leftarrow\phi$
3.4 Algorithms
The two stages of training K-CROSS are depicted in Fig 2. The details of two-stage training algorithms and the inference algorithm are described in Algorithm 1, Algorithm 2 and Algorithm 3, respectively. For clarity, Table 1 provides notation descriptions that occurred in our algorithms.
4 NIRPS Dataset and Radiologist Score
To comprehensively evaluate the synthesis performance, we construct a large-scale multi-modal neuroimaging perceptual similarity (NIRPS) dataset with 6,000 radiologist judgments. NIRPS dataset is composed of three subsets generated by CycleGAN (Zhu et al., 2017), MUNIT Huang et al. (2018) and UNIT Liu et al. (2017a). Each set contains 800 images generated by IXI and 1,200 images generated by BraTS. The IXI dataset includes two modalities, PD and T2, while the BraTS dataset includes three modalities, T1, T2, and FLAIR. In both the IXI and BraTS datasets, we randomly select 10 slices for training and collect the training results after each epoch of the model trained over 40 epochs.
IXI (Aljabar et al., 2011) collects nearly 600 MR images from normal and healthy subjects in three hospitals. The MR image acquisition protocol for each subject includes T1, T2, PD-weighted images (PD), MRA images, and Diffusion-weighted images. In this paper, we only use T1 (581 cases), T2 (578 cases) and PD (578 cases) data to conduct our experiments, and select the paired data with the same ID from the three modes. The image has a non-uniform length on the z-axis with the size of $256$ on the x-axis and y-axis. The IXI dataset is not divided into a training set and a test set. Therefore, we randomly split the whole data as the training set (80%) and the test set (20%).
BraTS2021 (Siegel et al., 2019; Bakas et al., 2017) is designed for the analysis and diagnosis of brain diseases. The dataset of multi-institutional and pre-operative MRI sequences is made publicly available and includes both training data (1251 cases) and validation data (219 cases). Each 3D volume is 155$\times$240$\times$240 in size and is imaged by four sequences: T1, T2, T1ce, and FLAIR.
Training Data Processing To ensure the validity and diversity of the data, we remove their skulls for each slice, by splitting the three-dimensional volume and choosing slices ranging from 50 to 80 on the z-axis. All images are cropped to $256$ pixels in size. During the training stage, we choose a total of 10k images from the IXI and BraTS2021 datasets.
@@ 4 NIRPS Dataset and Radiologist Score
4.0.1 Radiologist Score
The NIRPS dataset contains radiologist scores (RS) resulting from manual annotation for each image. It is worth noting that the radiologist score $RS$ includes 10 levels, i.e., $RS\in[0,0.1,0.2,..,0.9]$. The higher RS value indicates better-synthesized neuroimage quality. Radiologists give scores according to the level of diagnosis and therapy using the synthesized neuroimage. Fig. 8 gives the distribution result of $RS$. We can see that synthesized performance varies among the three models and the average $RS$ is in the middle.
4.0.2 How Radiologists Assess?
@@
4.0.3 How Radiologists Combine Their Evaluations?
We hire 10 radiologists to evaluate the quality of each synthesized neuroimage. We remove the highest score and the lowest score from all radiologists. The final score is then averaged by the rest score from 8 radiologists.
@@ 4.0.3 How Radiologists Combine Their Evaluations?
5 Experiment and Ablation Study
Table 2: Performance comparison of various IQA methods on NIRPS. NIRPS includes three subsets generated by CycleGAN, MUNIT, and UNIT. Specifically, each set consists of 800 images for IXI and 1,200 images for BraTS. Here, IXI and BraTS training sets are used to train our model. The inconsistency is computed, where $\downarrow$ is better. The best result is highlighted in red.
@@ 5 Experiment and Ablation Study
MAE (Chai and Draxler, 2014) 0.3127 0.3335 0.1807 0.2368 0.2462 0.2462
PSNR (Huynh-Thu and Ghanbari, 2008) 0.3387 0.2060 0.1832 0.2515 0.2387 0.2585
SSIM (Wang et al., 2004) 0.3333 0.1998 0.1957 0.2437 0.2523 0.2645
MS-SSIM (Wang et al., 2003) 0.3245 0.1973 0.1934 0.2425 0.2487 0.2521
NLPD (Laparra et al., 2016) 0.3128 0.1961 0.1921 0.2408 0.2466 0.2445
GMSD (Xue et al., 2014) 0.3069 0.1952 0.1910 0.2395 0.2421 0.2432
DeepIQA (Bosse et al., 2018) 0.3052 0.1943 0.1827 0.2380 0.2402 0.2410
LPIP (Zhang et al., 2018) 0.3023 0.1921 0.1786 0.2285 0.2398 0.2327
DIST (Ding et al., 2022) 0.3012 0.1901 0.1725 0.2291 0.2387 0.2230
@@
5.1 K-CROSS vs Other Metrics
Table 2 illustrates the inconsistency between metrics and human evaluations of several datasets and generative models, with the highest performance shown in red. The calculation method for inconsistency value is given in Algorithm 4. We evaluate K-CROSS on datasets created by CycleGAN, MUNIT, and UNIT. The first column indicates various IQA methods. The second column indicates which datasets were used to train the K-CROSS model, including IXI or BraTS. From Table 2, our proposed K-CROSS is more compatible with the assessments of radiologists. Note that the IXI dataset is a healthy person dataset. There is no lesion for each neuroimage. So K-CROSS only use the tumor branch and complex branch to assess the quality of neuroimage. The details are described in Section 5.3.
5.2 K-Space Importance
Table 3 records the ablation study of individual branches (the complex branch, tumor branch and structure branch) on various datasets. For example, when we conduct the ablation study of the complex branch, K-CROSS removes the tumor branch and the structure branch in the inference phase. In other words, K-CROSS only obtains $\eta_{complex}$ score. It applies the same setting for the other branches. It can be clearly observed that the complex branch obtains the highest score among the three branches. It strongly indicates the importance of k-space, which reflects the inherent properties of magnetic resonance imaging principles. The second best is the tumor branch. It also verifies the effectiveness of the tumor branch for the lesion disease dataset.
Table 3: Effect of k-space importance (complex branch) and tumor branch. The inconsistency is computed, where $\downarrow$ is better. The best result is in red and the second best result is in blue.
@@ 5.2 K-Space Importance
5.3 Metrics for Healthy Person
Table 4 shows K-CROSS performance that surpasses the mainstream IQA methods on the IXI healthy-person dataset. As for assessing the synthesised neuroimage of healthy persons, K-CROSS removes the tumor branch in the inference phase. Because there are no lesions on healthy person datasets. It means that K-CROSS only combines $\eta_{complex}$ and the score of the structure encoder as the final score. From Table 4, it can be obviously observed that K-CROSS complex branch score $\eta_{complex}$ (blue value) has surpassed the other IQA methods, which identify the importance of k-space for MRI of healthy persons. Thus, K-CROSS still be able to serve as the metric for the synthesized quality of healthy person’s neuroimage.
Table 4: Metric performance for healthy persons on IXI. The inconsistency is computed, where $\downarrow$ is better. The best result is in red and the second best result is in blue.
@@ 5.3 Metrics for Healthy Person
MAE (Chai and Draxler, 2014) 0.3160 0.1809 0.2442
PSNR (Huynh-Thu and Ghanbari, 2008) 0.3388 0.1835 0.2419
SSIM (Wang et al., 2004) 0.3352 0.1962 0.2423
MS-SSIM (Wang et al., 2003) 0.3238 0.1945 0.2418
NLPD (Laparra et al., 2016) 0.3135 0.1918 0.2409
GMSD (Xue et al., 2014) 0.3090 0.1908 0.2398
DeepIQA (Bosse et al., 2018) 0.3043 0.1856 0.2381
LPIP (Zhang et al., 2018) 0.3022 0.1810 0.2374
DIST (Ding et al., 2022) 0.3056 0.1805 0.2356
@@
5.4 Segmentation Network Effect
Table 5 shows K-CROSS remains stable performance even using different state-of-the-art medical segmentation models. Note that the parameters of the pre-trained segmentation network are frozen during the training phase. The first column denotes the segmentation method. We calculate the variance score of the K-CROSS value for CycleGAN, MUNIT and UNIT by using different segmentation backbone models. We find that the variance of K-CROSS performance is tiny (0.2%, 0.3%, and 0.2%). Hence, the performance of K-CROSS is not affected by the segmentation model.
Table 5: The effect of segmentation network for K-CROSS value for various generative models of NIRPS.
@@ 5.4 Segmentation Network Effect
nnUnet (Isensee et al., 2020) 0.1872 0.2246 0.2160
AttnUnet (Schlemper et al., 2019) 0.1873 0.2268 0.2157
SETR (Zheng et al., 2021) 0.1852 0.2254 0.2162
CoTr (Xie et al., 2021) 0.1856 0.2244 0.2154
TransUNet (Chen et al., 2021) 0.1862 0.2247 0.2162
SwinUNet (Cao et al., 2021) 0.1845 0.2252 0.2158
Table 6: K-CROSS on NIRPS (CycleGAN). The inconsistency is computed, where $\downarrow$ is better. The best result is in red and the second best result is in blue.
@@
MAE (Chai and Draxler, 2014) 0.3208 0.3321 0.3228
PSNR (Huynh-Thu and Ghanbari, 2008) 0.3190 0.3270 0.3261
SSIM (Wang et al., 2004) 0.3231 0.3231 0.3212
MS-SSIM (Wang et al., 2003) 0.3189 0.3180 0.3128
NLPD (Laparra et al., 2016) 0.3176 0.3163 0.3090
GMSD (Xue et al., 2014) 0.3067 0.3154 0.3067
DeepIQA (Bosse et al., 2018) 0.3058 0.3124 0.3045
LPIP (Zhang et al., 2018) 0.3033 0.3056 0.3039
DIST (Ding et al., 2022) 0.3021 0.3032 0.3006
@@
Table 7: K-CROSS on the NIRPS dataset (MUNIT). The inconsistency is computed, where $\downarrow$ is better. The best result is in red and the second best result is in blue.
@@
MAE (Chai and Draxler, 2014) 0.1872 0.1865 0.1842
PSNR (Huynh-Thu and Ghanbari, 2008) 0.1853 0.1873 0.1853
SSIM (Wang et al., 2004) 0.1967 0.1986 0.1982
MS-SSIM (Wang et al., 2003) 0.1932 0.1952 0.1923
NLPD (Laparra et al., 2016) 0.1945 0.1943 0.1921
GMSD (Xue et al., 2014) 0.1920 0.1934 0.1932
DeepIQA (Bosse et al., 2018) 0.1835 0.1845 0.1856
LPIP (Zhang et al., 2018) 0.1783 0.1754 0.1784
DIST (Ding et al., 2022) 0.1742 0.1731 0.1786
@@
Table 8: K-CROSS on the NIRPS dataset (UNIT). The inconsistency is computed, where $\downarrow$ is better. The best result is in red and the second best result is in blue.
@@
MAE (Chai and Draxler, 2014) 0.2468 0.2445 0.2480
PSNR (Huynh-Thu and Ghanbari, 2008) 0.2421 0.2410 0.2410
SSIM (Wang et al., 2004) 0.2545 0.2505 0.2470
MS-SSIM (Wang et al., 2003) 0.2486 0.2474 0.2463
NLPD (Laparra et al., 2016) 0.2481 0.2461 0.2450
GMSD (Xue et al., 2014) 0.2410 0.2388 0.2478
DeepIQA (Bosse et al., 2018) 0.2398 0.2376 0.2391
LPIP (Zhang et al., 2018) 0.2352 0.2341 0.2384
DIST (Ding et al., 2022) 0.2348 0.2337 0.2386
@@
5.5 General Metric? Overcoming Domain Gap
The purpose of this paper is to demonstrate that K-CROSS is capable of serving as the standard measure for MRI datasets. We conduct extensive experiments and the results are given in Table 6, Table 7, and Table 8 to demonstrate that K-CROSS is not affected by dataset domain gap and the generative model. The training dataset is the BraTS dataset, and the test dataset is IXI. As described in Section 5.3, we remove the tumor branch score, when K-CROSS evaluates the quality of neuroimage in healthy cases. We also observe that K-CROSS averagely surpasses DIST and LPIP (SOTA for natural image) by 7.8% and 16.5%, respectively, which proves that K-CROSS is built upon the basis of MRI principle instead of only on the natural image level. From this ablation study, we demonstrate that the performance of K-CROSS ($\mathcal{L}_{stru}$+$\mathcal{L}_{freq}$) is stable across several MRI datasets, with the potential to serve as a generic measure for evaluating the quality of the synthesized MRI.
6 Conclusion
In this paper, we proposed a new metric K-CROSS to evaluate the performance of medical images synthesized, which is based on the principle of magnetic resonance imaging. To improve the reconstruction capability during K-CROSS training, a complex U-Net was developed. As for training a learning-based full IQA metric, we further constructed a large-scale multi-modal neuroimaging perceptual similarity (NIRPS) dataset. The experimental results indicate that K-CROSS is a useful indicator for evaluating the quality of the medical data generated. However, our method heavily relies on deep learning-based techniques but without directly injecting the knowledge of radiologists into K-CROSS. In the future, K-CROSS should combine causal inference methods to improve interpretability.